OpenStudy (nuccioreggie):

Two functions, A and B, are described as follows: Function A y = 8x + 3 Function B The rate of change is 1 and the y-intercept is 4. How much more is the rate of change of function A than the rate of change of function B? 1 7 8 9

OpenStudy (nuccioreggie):

@amyhashimoto77

OpenStudy (amyhashimoto77):

The rate of change is the slope. Function A slope: 8 Function B slope: 1 [Equation: y = x + 4]

OpenStudy (amyhashimoto77):

The question confuses me because it say "How much more?" and usually in math they mean add. So 7 more. But if its asking, "how many TIMES more" than the answer would be 8

OpenStudy (nuccioreggie):

so its B?

OpenStudy (amyhashimoto77):

Well, if this is a one chance only, than I would go with 8.

OpenStudy (nuccioreggie):

The length of a rectangle is 7 units and its width is 5 units. What is the approximate length of the diagonal of the rectangle? 4.9 units 6 units 8.6 units 9 units

OpenStudy (amyhashimoto77):

C

OpenStudy (amyhashimoto77):

In this one, you split the rectangle diagonally and you end up with 2 triangles. And then use the pythagorean theorem. 7^2 + 5^2 = c^2 49 + 25 = c^2 74 =c^2 √74 = c 8.6 = c

OpenStudy (nuccioreggie):

A cylindrical well is 15 meters deep and has a diameter of 1.6 meters. Approximately how many cubic meters of soil were dug out to make the well? (Use π = 3.14.) 30.14 cubic meters 37.68 cubic meters 120.57 cubic meters 188.40 cubic meters

OpenStudy (nuccioreggie):

i think its D but im not sure

OpenStudy (amyhashimoto77):

You need to find the volume of a cylinder, which is (pi)r^2 x h The radius is half the diameter. r = 1.6/2 = 0.8 The height is depth of the well. h = 15 Now use the equation: (3.14)(0.8^2) x (15) = 30.14 m^3

OpenStudy (nuccioreggie):

ohhhh ok so that would be A right oh ok

OpenStudy (nuccioreggie):

In triangle PQR, r2 = p2 + q2. Triangle PQR has sides p, q, r opposite to the corresponding vertices P, Q, R Which equation is true about the measure of the angles of the triangle? The measure of angle PRQ is equal to 100 degrees The measure of angle PRQ is equal to 90 degrees The measure of angle PQR plus the measure of angle QPR is equal to 80 degrees The measure of angle PQR plus the measure of angle QPR is equal to 70 degrees

OpenStudy (amyhashimoto77):

B

OpenStudy (nuccioreggie):

Two equations are shown: Equation A y = −3x − 2 Equation B y equals 3 over x plus 5 Which statement best compares the graphs of the two equations? Both are nonlinear. Both are linear. Equation A is nonlinear and Equation B is linear. Equation A is linear and Equation B is nonlinear.

OpenStudy (amyhashimoto77):

D

OpenStudy (nuccioreggie):

can you explain how itsD?

OpenStudy (nuccioreggie):

Orio and Ted buy an equal number of comic books every month. The following equation shows the number of comic books, y, that Orio has after x months: y = 2x + 15 The following graph shows the number of comic books, y, that Ted has after x months: Graph shows Months on x axis and Number of Comic Books on y axis. The scale on x axis is shown from 0 to 5 at increments of 1. The scale on y axis is shown from 0 to 14 at increments of 2. The title of the graph is Teds Comic Books. A straight line is shown which joins the ordered pair 0, 6 and 4, 14 After 2 months, how many more comic books does Orio have than Ted? 1 2 4 9

OpenStudy (amyhashimoto77):

Equation A is linear because it's x to the first power Linear = straight line Equation B, I'm not too sure how to explain, but I plugged y=(3/x) + 5 into a graphing calculator and got something like this... hold on And so Equation B not linear.

OpenStudy (nuccioreggie):

OK THANK YOU FOR YOUR EXAMPLE (:

OpenStudy (nuccioreggie):

Orio and Ted buy an equal number of comic books every month. The following equation shows the number of comic books, y, that Orio has after x months: y = 2x + 15 The following graph shows the number of comic books, y, that Ted has after x months: Graph shows Months on x axis and Number of Comic Books on y axis. The scale on x axis is shown from 0 to 5 at increments of 1. The scale on y axis is shown from 0 to 14 at increments of 2. The title of the graph is Teds Comic Books. A straight line is shown which joins the ordered pair 0, 6 and 4, 14 After 2 months, how many more comic books does Orio have than Ted? 1 2 4 9

OpenStudy (nuccioreggie):

THE ANSWER CHOICE IS b?

OpenStudy (anonymous):

C

OpenStudy (nuccioreggie):

@amyhashimoto77 is he correct?

OpenStudy (amyhashimoto77):

Haha I got D

OpenStudy (nuccioreggie):

ok lol

OpenStudy (amyhashimoto77):

Orio: y=2x + 15 Ted: y=2x + 6 y=2(2)+15 y=4+15 y=19 Orio has 19 comics in 2 months y=2x+6 y=2(2)+6 y=4+6 y=10 Ted has 10 comics in 2 months SUBTRACT: 19 - 10 = 9!!!